OpenStudy (anonymous):

The sixth term of a sequence is 128 and the common ratio is 2. Use the explicit formula for a geometric sequence to find the first term.

OpenStudy (shamim):

given t6=128 r=2

OpenStudy (shamim):

we know the formula is tn=ar^(n-1)

OpenStudy (shamim):

so we can write t6=ar^(6-1)

OpenStudy (shamim):

128=a*2^(6-1)

OpenStudy (shamim):

tell me a=?